Barium hydroxide crystallises as one of several possible hydrates. A sample of barium hydroxide is analysed in order to determine its formula. 3.632 g of the sample was dissolved in water to make 250.0 mL of solution. 25.00 mL of this solution was titrated with 0.0987 M HCl, using methyl orange indicator. Titres of 23.26, 23.29 and 23.34 mL were obtained.

Determine the formula (that is, find the value of x) of the barium hydroxide sample.
1
Expert's answer
2020-09-08T05:57:57-0400

Step I : Get the average titre of the acid

"\dfrac{23.26+23.29+23.34}{3} =\dfrac{69.89}{3}=23.29cm3"


Step II : Calculate moles of HCl in this volume


We know that, concentration of the acid is 0.0987M


"\therefore 0.0987moles =1000cm3"


"? =23.29cm"


"=\dfrac{0.0987 x 23.29}{1000}"


"=0.0022999mol"

These are the moles of HCl that reacted with barium hydroxide solution.


Step III : Write balanced equation for the reaction and calculate moles of barium hydroxide that reacted.

Ba(OH)2(aq) + 2HCl(aq) "\to" BaCl2(aq) + 2H2O(l)


From the reaction mole ratio (Moles of base:moles of acid) is 1:2

Meaning 1 mole of Ba(OH)2 reacts with 2 moles of HCl


"\therefore moles of barium hydroxide= \dfrac{1}{2} x moles of HCl""Moles of barium hydroxide = \dfrac{1}{2}x0.0022999=0.00114969 moles"


These moles of Ba(OH)2 were in 25cm3

How many moles were there in the 250cm3 of the stock solution?

By cross multiplication we get;

"=\dfrac{0.00114969 x 250}{25} = 0.0114969 moles"


These were the number of moles in the 3.632g of solid sample that was dissolved.


Step IV : Calculate RFM of the solid sample of Ba(OH)2 . xH2O


From the formula; "Moles = \dfrac{mass}{RFM}" we get,


"RFM = \dfrac{mass}{moles}"


"RFM = \dfrac{3.632}{0.0114969} = 315.9111083 amu"


Step V : Now Determine the value of x in Ba(OH)2. xH2O


(Atomic mass, Ba=137.3, H=1.0, O=16.0)


Ba(OH)2.xH2O = 315.911083


173.3 + 2(16+1) + x(2+16) = 315.911083

137.3 + 34 + 18x = 315.911083


18x = 315.911083 - 171.3

18x = 144.61

Divide both sides by 18

"x=\dfrac{144.61}{18} = 8.034"

The value of x is approximately 8

The formula of hydrated barium hydroxide is therefore Ba(OH)2 . 8H2O